The number of telephone calls that arrive at a phone exchange is often modeled as a Poisson random variable. Assume that on the average there are λ= 10 calls per hour. Determine the following probabilities:


1
Expert's answer
2021-11-19T15:14:51-0500

for Poisson distribution:


"P(x=k)=\frac{\lambda^ke^{-\lambda}}{k!}"


for example, probability to get 2 calls per hour:


"P(x=2)=\frac{10^2e^{-10}}{2!}=0.0023"